# Three Quenched, Faint Dwarf Galaxies in the Direction of NGC 300: New Probes of Reionization and Internal Feedback

### v1

@@ 1 Introduction
To complement and extend observations around Milky Way-mass systems, and to test the astrophysics input into modern simulations, it is necessary to identify and understand the lowest-mass dwarf galaxies in a variety of environments. One area of interest is the population of dwarfs in lower density environments: at the edge of the Local Group’s main galaxies (e.g. McQuinn et al., 2023, 2024), around less massive hosts such as Magellanic Cloud-like systems (e.g. Sand et al., 2015a; Carlin et al., 2016, 2021; McNanna et al., 2024; Davis et al., 2024), and in isolation (e.g. Cannon et al., 2011; McQuinn et al., 2014, 2021; Sand et al., 2022; Jones et al., 2023, 2024; Li et al., 2024). The effects of tidal and ram pressure stripping will be diminished in low density environments (e.g. Garling et al., 2024), with the exception of ‘backsplash’ systems (e.g. Teyssier et al. 2012; Buck et al. 2019; Santos-Santos et al. 2023), and so offer an opportunity to cleanly study other quenching mechanisms, such as reionization (Bullock et al., 2000; Benson et al., 2002; Ricotti & Gnedin, 2005; Jeon et al., 2017; Applebaum et al., 2021) and star formation/supernova feedback (e.g. Dekel & Silk, 1986; Mac Low & Ferrara, 1999; Hopkins et al., 2012; El-Badry et al., 2018).
Here we report the discovery of three faint, seemingly quenched, dwarf galaxies in the direction of NGC 300. NGC 300 is an SA spiral galaxy at D=2.0 Mpc (based on the tip of the red giant branch; Dalcanton et al., 2009) in the direction of the Sculptor group. It has a K-band luminosity very similar to the Large Magellanic Cloud (Mutlu-Pakdil et al., 2021), making it an excellent comparison for satellite population studies. Throughout this work, we assume that NGC 300 has a virial radius of 120 kpc and halo mass of log(/)11.29 as calculated in Mutlu-Pakdil et al. (2021), using the stellar mass–halo mass relation of Moster et al. (2010). These values are uncertain, but are used to help ascertain the relationship between the three faint dwarf discoveries and NGC 300, with this caveat in mind. We present the discovery of the three dwarf galaxies in Section 2, and discuss follow-up optical observations in Section 3. In Section 4 we measure the basic physical properties of the dwarfs (distance, stellar population, gas content, structure and luminosity). We discuss the environment of each in Section 5 and compare them with other dwarfs in low density environments. Finally, we summarize and conclude in Section 6.
@@ 2 Ultra-faint Dwarf Galaxy Discovery Around NGC 300
Following the recent discovery of Tucana B through a focused visual search utilizing the DESI Legacy Imaging Surveys and its interactive color image viewer111https://www.legacysurvey.org/viewer (Sand et al., 2022), we began a systematic search for faint companions around other nearby galaxies (D2 Mpc), including NGC 300 (for reference, Data Release 9 of the DESI Legacy Imaging Surveys was used). We uploaded a custom file to mark off a region with a projected box size of 400 kpc on a side centered on NGC 300 itself (11.3 deg), and visually searched for stellar over-densities with underlying diffuse light. This coverage area out to 200 kpc encloses the full virial radius of NGC 300, which is approximately 120 kpc (Mutlu-Pakdil et al., 2021). Faint dwarf galaxies at the distance of NGC 300 can resemble ‘semi-resolved’ objects, with both resolved and diffuse components, at the depth of DECaLS (23.5-24 mag; depending on the field; Dey et al. 2019). We searched the NGC 300 footprint several times, at different spatial scales and contrast levels.
Three high-confidence faint dwarf galaxy candidates stood out during the search, and we show color cutouts obtained from the Legacy Survey Image Viewer in Figure 1. We also present a schematic of the footprint searched, and the projected position of these discoveries in Figure 2. NGC 300 and surrounding areas are in the footprint of the Dark Energy Survey Data Release 2 (DES DR2; Abbott et al., 2021), and we downloaded photometry of the field using NOIRLab’s Query Interface Tool222https://datalab.noirlab.edu/query.php. This photometry was suggestive of an old, metal-poor red giant branch (RGB) at 2 Mpc but was ultimately inconclusive. We thus sought deeper ground-based optical imaging, which we discuss next.
@@ 4.1 Distance
For consistency, we thus measure the distance to the three dwarf galaxies using a CMD-fitting technique, comparing the number of stars consistent with an old, metal-poor theoretical isochrone as a function of assumed distance. A similar technique has been used to measure the distance to many Milky Way ultra-faint dwarf galaxies (e.g. Walsh et al., 2008; Sand et al., 2009), as well as Tucana B (Sand et al., 2022). In our analysis we use Dartmouth isochrones (Dotter et al., 2008) with a 13.5 Gyr stellar population and low metallicities ([Fe/H]=2.5, 2.0 and 1.5.). For each dwarf we include all stars brighter than the 50% completeness threshold, and stars spatially coincident with the main body of each dwarf (a radius of 0.3–0.5′), assessed visually. Then, we shift each isochrone fiducial through distance moduli () between 25.0 and 28.0 mag (1.0 and 4.0 Mpc) in 0.025 mag steps, counting the number of stars consistent with the isochrone at that distance. A scaled background region at the edge of the GMOS field of view is used to subtract off contaminants for each distance modulus trial. For a given , a star is considered consistent with the isochrone if its color is within a red/blue boundary derived from our photometric uncertainties. The best distance corresponds to the distance modulus which maximizes the number of stars consistent with the isochrone after background removal.
The fits to the [Fe/H]=2.0 isochrone are excellent, as are those for the [Fe/H]=2.5 isochrone, while the best-fitting [Fe/H]=1.5 isochrone are slightly too red and do not capture the shape of the observed RGB. For this reason, we adopt the [Fe/H]=2.0 distance in this work (see the second-left panel in Figure 3), with the range of best-fitting distance values for the full metallicity set as our uncertainty. We list these distances in Table 1.
@@ 4.4 Gas Content
We downloaded GASS spectral line cubes333https://www.astro.uni-bonn.de/hisurvey/gass/index.php for each dwarf and extracted a spectrum over spatial pixels, centered on each target. We inspected these spectra (and the cubes) at spectral resolutions of 5 and 25 km s-1, but saw no significant emission features other than the MW and high-velocity clouds. We measured the rms noise at 25 km s-1 resolution for each dwarf and assuming 25 km s-1 as a conservative fiducial velocity width, estimated the upper limits in each case. These limits are given in Table 1. More stringent gas limits will be necessary to verify that these systems are truly gas free (down to /LV1; e.g. Putman et al., 2021).

### v2

@@ 1 Introduction
To complement and extend observations around Milky Way-mass systems, and to test the astrophysics input into modern simulations, it is necessary to identify and understand the lowest-mass dwarf galaxies in a variety of environments. One area of interest is the population of dwarfs in lower density environments: at the edge of the Local Group’s main galaxies (e.g. McQuinn et al., 2023, 2024), around less massive hosts such as Magellanic Cloud-like systems (e.g. Sand et al., 2015a; Carlin et al., 2016, 2021, 2024; McNanna et al., 2024; Davis et al., 2024), and in isolation (e.g. Cannon et al., 2011; McQuinn et al., 2014, 2021; Sand et al., 2022; Jones et al., 2023, 2024; Li et al., 2024). The effects of tidal and ram pressure stripping will be diminished in low density environments (e.g. Garling et al., 2024), with the exception of ‘backsplash’ systems (e.g. Teyssier et al. 2012; Buck et al. 2019; Santos-Santos et al. 2023), and so offer an opportunity to cleanly study other quenching mechanisms, such as reionization (Bullock et al., 2000; Benson et al., 2002; Ricotti & Gnedin, 2005; Jeon et al., 2017; Applebaum et al., 2021) and star formation/supernova feedback (e.g. Dekel & Silk, 1986; Mac Low & Ferrara, 1999; Hopkins et al., 2012; El-Badry et al., 2018).
Here we report the discovery of three faint, seemingly quenched, dwarf galaxies in the direction of NGC 300. NGC 300 is an SA spiral galaxy at D=2.0 Mpc (based on the tip of the red giant branch; Dalcanton et al., 2009) in the direction of the Sculptor group. It has a K-band luminosity very similar to the Large Magellanic Cloud (Mutlu-Pakdil et al., 2021), making it an excellent comparison for satellite population studies111If we use a -band mass-to-light ratio of 1, both NGC 300 and the LMC have a stellar mass of 2.6109 . This is consistent with stellar mass estimates of 2109 for NGC 300 (Muñoz-Mateos et al., 2015) and 2.7109 for the LMC (van der Marel et al., 2002) in the literature.. Throughout this work, we assume that NGC 300 has a virial radius of 120 kpc and halo mass of log(/)11.29 as calculated in Mutlu-Pakdil et al. (2021), using the stellar mass–halo mass relation of Moster et al. (2010). These values are uncertain, but are used to help ascertain the relationship between the three faint dwarf discoveries and NGC 300, with this caveat in mind. We present the discovery of the three dwarf galaxies in Section 2, and discuss follow-up optical observations in Section 3. In Section 4 we measure the basic physical properties of the dwarfs (distance, stellar population, gas content, structure and luminosity). We discuss the environment of each in Section 5 and compare them with other dwarfs in low density environments. Finally, we summarize and conclude in Section 6.
@@ 2 Ultra-faint Dwarf Galaxy Discovery Around NGC 300
Following the recent discovery of Tucana B through a focused visual search utilizing the DESI Legacy Imaging Surveys and its interactive color image viewer222https://www.legacysurvey.org/viewer (Sand et al., 2022), we began a systematic search for faint companions around other nearby galaxies (D2 Mpc), including NGC 300 (for reference, Data Release 9 of the DESI Legacy Imaging Surveys was used). We uploaded a custom file to mark off a region with a projected box size of 400 kpc on a side centered on NGC 300 itself (11.3 deg), and visually searched for stellar over-densities with underlying diffuse light. This coverage area out to 200 kpc encloses the full virial radius of NGC 300, which is approximately 120 kpc (Mutlu-Pakdil et al., 2021). Faint dwarf galaxies at the distance of NGC 300 can resemble ‘semi-resolved’ objects, with both resolved and diffuse components, at the depth of DECaLS (23.5-24 mag; depending on the field; Dey et al. 2019). We searched the NGC 300 footprint several times, at different spatial scales and contrast levels.
Three high-confidence faint dwarf galaxy candidates stood out during the search, and we show color cutouts obtained from the Legacy Survey Image Viewer in Figure 1. We also present a schematic of the footprint searched, and the projected position of these discoveries in Figure 2. NGC 300 and surrounding areas are in the footprint of the Dark Energy Survey Data Release 2 (DES DR2; Abbott et al., 2021), and we downloaded photometry of the field using NOIRLab’s Query Interface Tool333https://datalab.noirlab.edu/query.php. This photometry was suggestive of an old, metal-poor red giant branch (RGB) at 2 Mpc but was ultimately inconclusive. We thus sought deeper ground-based optical imaging, which we discuss next.
@@ 4.1 Distance
For consistency, we thus measure the distance to the three dwarf galaxies using a CMD-fitting technique, comparing the number of stars consistent with an old, metal-poor theoretical isochrone as a function of assumed distance. A similar technique has been used to measure the distance to many Milky Way ultra-faint dwarf galaxies (e.g. Walsh et al., 2008; Sand et al., 2009), as well as Tucana B (Sand et al., 2022). In our analysis we use Dartmouth isochrones (Dotter et al., 2008) with a 13.5 Gyr stellar population and low metallicities ([Fe/H]=2.5, 2.0 and 1.5.). For each dwarf we include all stars brighter than the 50% completeness threshold, and stars spatially coincident with the main body of each dwarf (a radius of 0.3–0.5′), assessed visually. Then, we shift each isochrone fiducial through distance moduli () between 25.0 and 28.0 mag (1.0 and 4.0 Mpc) in 0.025 mag steps, counting the number of stars consistent with the isochrone at that distance. A scaled background region at the edge of the GMOS field of view is used to subtract off contaminants for each distance modulus trial. For a given , a star is considered consistent with the isochrone if its color is within a red/blue boundary derived from our photometric uncertainties. The best distance corresponds to the distance modulus which maximizes the number of stars consistent with the isochrone after background removal. To assess our uncertainties on the distance measurement, we bootstrap (resample with replacement) both the input dwarf stars, and the background objects.
The fits to the [Fe/H]=2.0 isochrone are excellent, as are those for the [Fe/H]=2.5 isochrone, while the best-fitting [Fe/H]=1.5 isochrone are slightly too red and do not capture the shape of the observed RGB. For this reason, we adopt the [Fe/H]=2.0 distance in this work (see the second-left panel in Figure 3), with the range of best-fitting distance values for the full metallicity set as our formal uncertainty. We choose this metric for the reported uncertainty rather than the results of the bootstrap analysis, as the bootstrap uncertainties are sub-dominant in comparison to the distance spread between metallicities. We list these distances in Table 1.
@@ 4.4 Gas Content
We downloaded GASS spectral line cubes444https://www.astro.uni-bonn.de/hisurvey/gass/index.php for each dwarf and extracted a spectrum over spatial pixels, centered on each target. We inspected these spectra (and the cubes) at spectral resolutions of 5 and 25 km s-1, but saw no significant emission features other than the MW and high-velocity clouds. We measured the rms noise at 25 km s-1 resolution for each dwarf and assuming 25 km s-1 as a conservative fiducial velocity width, estimated the upper limits in each case. These limits are given in Table 1. More stringent gas limits will be necessary to verify that these systems are truly gas free (down to /LV1; e.g. Putman et al., 2021).